# Leveraging Self-Supervised Learning for Scene Recognition in Child Sexual Abuse Imagery

### v1

@@ III-A Problem Definition
Any study with CSAI must consider the ethical concerns surrounding such data. It is common then to always design methods under a protocol not dissimilar to the Domain Generalization (DG) task.
@@
The remaining challenge is then to define which $\mathcal{S}^{train}$ domains can be used to train such a robust model, as well as the training procedure itself. As presented previously, our chosen classification task is Indoor Scene Classification. Whereas scene recognition may be acknowledged as a homogeneous classification task [37], indoor and outdoor environments vary greatly in textures, objects, colors, shadows, framing, and many other features. Indoor environments contain large non-textured regions – rare in outdoor environments – making models that work well on outdoor images fail at indoor scenes [81]. Moreover, the distribution and variety of objects are much denser in indoor environments, making semantic segmentation more expensive and prone to error [63].
To solve Indoor Scene Classification with CSAI in mind, we propose to compare the performance of a classic transfer learning approach of fine-tuning a network pre-trained on ImageNet – our baseline – against a selection of self-supervised models known for generalizing to a plethora of downstream tasks – in some instances surpassing their supervised counterparts [29]. By comparing the results on our Indoor Scene Classification benchmark, we are able to select the best-performing model for this task, the code, and learned weights, which are then sent to law enforcement partners to do the final assessment on real CSAI.
@@ III-B Datasets
As mentioned in Section II-A, SSL is trained in two stages; the self-supervised stage trains a model under a pretext task, a pseudo-labeled classification and representation learning task meant to instill general representations into the model. This pre-trained model is then fine-tuned on a downstream task and evaluated. For each of these stages, (i) pretext, (ii) downstream, and (iii) evaluation, we have designed new sub-sets of existing public datasets, leaving (iv) the final CSAI evaluation and related dataset in the hands of partner law enforcement agents. Creating new sub-sets was necessary because most scene recognition datasets with enough samples comprise both outdoor and indoor classes, and we wish to filter out the former. We present the selection for each stage next.
@@
We propose two combined datasets for the pretext self-supervised task: Indoors.all and Indoors.real. The former comprises all chosen indoor images from Hypersim [82], OpenRooms [83], InteriorNet [84], and Places365-Challenge [39], while the latter is then simply the Places365-Challenge indoor images. With this in mind, Indoors.all is called so as it contains not only real but also synthetic images and their rendered views (depth, segmentation, etc.). These views are used as pseudo-labels in the SSL pretext task, which we believe is more informative than random augmentations and could potentially produce better models, inspired by Ren and Lee [63]. Table I summarizes the included datasets and their sizes.
TABLE I: Datasets used for the self-supervised pretext task. Size stands for the total number of images, and Synth stands for synthetic datasets. ${}^{\dagger}$ombinations of the above public datasets.
@@
Indoors.all (${}^{\dagger}$ 7,778,152
Indoors.real (${}^{\dagger}$ 2,490,632
While most SSL models are readily available pre-trained on the well-known ImageNet dataset, we corroborate the finding from Zhou et al. [39] that shows that models trained on object-centric datasets do not perform well on scene-centric tasks and vice-versa and even that training on both distributions may be beneficial. This comparison is further discussed with the experimental results in Section V.
@@
We introduce a new subset of Places – called Places8 – where classes are selected to highlight environments most common in CSAI. This is a smaller dataset than the ones used for the pretext task; it represents our downstream task and is used for fine-tuning the model post self-supervised learning.
@@
Class Test Train Val %
Original Categories
bathroom
5,740 51,655 200 13.4
bathroom, shower
bedroom
11,112 100,012 600 25.9
bedchamber, bedroom, hotel room, berth, dorm room, youth hostel
child’s room
4,650 41,849 300 10.8
child’s room, nursery, playroom
classroom
3,751 33,763 200 8.7
classroom, kindergarden classroom
dressing room
2,432 21,889 200 5.7
closet, dressing room
living room
9,940 89,458 500 28.7
home theater, living room, recreation room, television room, waiting room
studio
1,404 12,633 100 3.3
television studio
swimming pool
1,505 13,547 200 3.5
jacuzzi, swimming pool
Total
40,534 364,806 2300 100
@@
Thus, we design a small “custom dataset” from online images to check if the model performance holds outside of the controlled nature of Places8. The dataset comprises 10 images per class comprising the 8 Places8 classes: bathroom, bedroom, child’s room, classroom, dressing room, living room, studio, and swimming pool.
The OOD Scenes set is a sample of images taken from Google images, Bing images, and the Dollar Street dataset [85]. All images are free to share, modify, and use, including Dollar Street, licensed under CC-BY 4.0 Commercial.
@@
For reproducibility, we will make Places8 and OOD Scenes available at Zenodo upon request.
III-C Experimental Design
We aim to study the performance of self-supervised methods on our Indoor Scene Classification task and apply the best model on CSAI. Given that we have our pipeline steps of (ii) downstream task, (iii) evaluation benchmark, and (iv) evaluation on CSAI fixed by design, our analysis will target the changes that can be enacted on the (i) pretext task of SSL. Our goal is to use the Places8 dataset – created with law enforcement guidance – to approximate performance in CSAI. Because the sensitive data is restrictive in its access, we do not hope to use it to do more than one evaluation with a single model; this best model is then to be selected through $k$-fold cross-validation ($k=5$) on the validation set of Places8.
We henceforth consider the pretext task through the lenses of three research questions, the first two pertaining to the data used to pre-train the model and the last to the method:
  1. Q1
    Does performing SSL on an object-centric dataset help or hurt downstream performance? Following the findings from Zhou et al. [39] that showed the combination of training on scene and object-centric datasets was beneficial to both tasks, we consider how SSL on ImageNet (an object-centric dataset) impacts the downstream performance on Places8. This is in addition to our comparison against a supervised baseline model pre-trained on ImageNet and fine-tuned on Places8.
  2. Q2
  3. Q3
With these questions and the experiments that will answer each in mind, our analysis can be summarized by the model variants yielded from the combination of the choices enlisted through the questions. When comparing these models, we use Bayesian estimation – replacing the common t-test – for ranking and contrasting. Bayesian Estimation Supersedes the T-test (BEST) [86] offers a nuanced and comprehensive view of experimental results compared to traditional t-tests. In our particular case, using 5-fold cross-validation breaks the independence assumption of the traditional t-test, leading to unreliable results. Also, we measure the advantage of using SSL over supervised techniques, which is accurately done by Bayesian methods. Fig. 1 illustrates this experimental pipeline.
Figure 1: Methodology pipeline. Self-supervised methods are often used in a common two-stage training protocol: (1) the pretext task – or pre-training stage – and (2) the downstream task – or fine-tuning stage. The pretext task uses unlabeled data and runs the SSL technique, while the downstream task uses labeled data from our target task. Our downstream stage (and target task) is fixed, so we consider the pretext task through the lenses of the three research questions depicted (described in Section III).
@@ IV Model Selection
In this section, we go through the model selection process. We recall that experiments that directly use CSAI are limited by law enforcement agents’ availability and local hardware requirements, and are in general discouraged. For this reason, we perform model selection with data that approximates the problem, hence our creation of Places8 in partnership with agents (Section III-B). The following experiments then sought to best answer the three questions posed on Section III by using 5-fold cross-validation and 3×3$\times$3 × experiment repetitions, comparing model variants on Places8 while avoiding spurious results from training randomization.
@@ IV-A Experimental Setup
We conducted all experiments using 6 NVIDIA RTX GPUs A6000 with 48 GB each. We used PyTorch 1.10222https://pytorch.org as the deep learning framework with the Vision library for state-of-the-art self-supervised learning (VISSL) [^3] https://vissl.ai. ResNet-50 [87] was chosen as the backbone for all experiments. ImageNet pre-trained models are used directly from public “model zoos” while scene-centric SSL pre-trained models are trained from scratch or ImageNet-based weights; all models are later fine-tuned on Places8. The training times were optimized through mixed precision and activation checkpointing – allowing training to progress at 1.5 hours/epoch. The batch size was set to 1024 samples per GPU ($1024\times 6$), the LARS optimizer was used with learning rate scheduler of cosine half wave 3$\times$ restarts given the high batch size we chose, which follows the same training hyperparameters and guidelines from VISSL [88]. SSL pre-training was performed in 50 epochs, given our training budget, while fine-tuning was done on 100 epochs for Places8 with early stopping supported by the validation loss.
@@ IV-B SSL on Object-centric Data
Because training within an SSL protocol is expensive (see Section IV-A), we sought to limit our exploration of this first question to a single SSL method and observe only the effect of adding object-centric learning to the mix; SwAV was the chosen method for this section for it is the best performing on the ImageNet classification downstream task of the four SSL methods (75.3% vs. 73.2% from second place Barlow Twins); we therefore study the 5 model variants yielded by questions Q1 and Q2 to isolate the impact of training first on ImageNet (Q1). The results can be found in Table III. This evaluation shows that significant improvements can be made using the widely available object-centric data to complement training, even when the downstream task is scene-centric. Comparing all experiments on all folds, we could measure with 97% confidence an average improvement of 3.1 percentage points in balanced accuracy on Places8 validation when starting with ImageNet SSL vs. models trained with scene-centric SSL from scratch.
@@ IV-C SSL on Scene-centric Data
When considering the inclusion of synthetic data, we compare SSL training with indoors.real versus indoors.all and found a 3% advantage to not using additional data (with 89.8% confidence). We hypothesize that the available synthetic data – while abundant in the number of samples – is limited in its diversity, with a large number of views being derived from the same 3D environment.
@@
TABLE IV: Comparing the impact of the inclusion scene-centric – real and synthetic – datasets for pre-training and the SSL training protocol. Results are reported as accuracy on the downstream Indoor Scene Recognition task on the Places8 dataset.
@@ IV-D The Self-Supervised Learning Protocol
Our selected model is then comprised of the answers, following the Barlow Twins SSL protocol with the pipeline of object-centric SSL with ImageNet, scene-centric SSL with indoors.real, and fine-tuning on the Places8 downstream task. The final model was trained end-to-end with the full training set of Places8 (no folds) and evaluated on its test set. The results showed then 71.6% balanced accuracy. This is higher than what is commonly shown for the full Places dataset [37], but with fewer classes, it is impossible to do a fair comparison. We hope future studies consider the use of Places8 given its relationship to the CSAI final task.
@@ V Results on OOD Data and CSAI
Finally, in this section, we test our validated model – fine-tuned with the full training set of Places8 – on our two evaluation datasets. First, we consider the newly introduced OOD Scenes (Section III-B), where images were selected to represent a sample distribution distinct from the Places dataset.
Through our partnership with law enforcement agents, we test how our Indoor Scene Classification model performs on real CSAI data for our second and final evaluation. This final experiment on CSAI was executed in a GPU workstation owned by and located on the partner agency. No training is performed with either set.
@@ V-B CSAI
These experiments are done through our partnership with Brazilian Federal Police expert agents; our validated model weights and inference code were sent to the agents who proceeded with running and collecting metrics.
Because Indoor Scene Recognition is a new task for CSAI data, the agents first selected 615 random samples from their allotted-for-research dataset. These samples were then annotated with scene labels. The randomness of the process showed that while a significant amount of the images were taken indoors (392), some were outdoors (118) or did not depict enough of the environment to allow for a label (105). The random selection showing most photos taken indoors corroborates the need for research into indoor scenes for CSAI. We follow with the 392 indoor scenes for our experiments.
Moreover, this subset is grouped into two categories of child abuse material: CSAI and Suspected CSAI. The CSAI category comprises the most serious and sensitive cases whilst Suspected CSAI does not show any nudity yet contains hints of eroticism. Among these, 313 images are from the CSAI category and 302 from the Suspected CSAI category. We show the distribution of samples between classes and between these two CSAI categories in Fig. 4. Classes have different frequencies per CSAI category, with “studio” being especially prevalent in Suspected CSAI (the agent affirms there is a correlation between this kind of scene and erotic posing). In contrast, the scenes of “classroom” and “dressing room” are not seen in any image. While this is partly due to the random selection process, it hints at a lower prevalence of these scenes, at least in the analyzed data.
@@ VI Discussion
We have shown that combining publicly available self-supervision models with large scene recognition datasets for a pretext task helps indoor classification on scenes prevalent in CSAI (but including synthetic data does not). The evaluation with real CSAI has, however, highlighted that these datasets are not representative of scenes recurrent in these sensitive images.
In particular, the ambiguity among classes such as “bedroom”, “child’s room”, and “living room” seems to be hard for the model to distinguish, particularly in CSAI, an ambiguity corroborated by the presence of correlation shift [90] between the Places and CSAI data distributions. The shift is evident in a few examples: the model correlates the presence of people in the “dressing room” class and the presence of children in “child’s room”, but in CSAI, there are people in all images (by definition) and most of those people are children regardless of the place depicted.
This work applied self-supervision to indoor classification to aid CSAI triage and recognition, showing that SSL can leverage scene images to perform better than supervised learning. SSL performance was compared in multiple conditions, using real and synthetic images, public and sensitive data. The inference in real CSAI is vital, exposing the real distribution of scenes in this data and how a model trained with publicly available data performs. The results obtained should serve as a starting point for further investigations.
@@ VII Challenges and Future Work
The current solution can classify at most two of the scene classes with acceptable performance, which can be used for filtering specifically for bathrooms and bedrooms but not much else. Our analysis is, of course, limited to Indoor Scenes, one aspect of CSAI, and we know a general solution must combine this with other aspects such as object and body part detection and people; the – expected – appearance of the latter was shown to particularly influence the performance of a model trained on samples with few people depicted. Given our findings regarding the importance of scenery to CSAI, corroborated during the data labeling by the agents, we hope our first step to be an incentive for more research to be done within the Indoor Scene Recognition field.
Additionally, all experiments used ResNet-50, which is fairly popular in SSL research, but multiple other architectures have gained much research interest after their sustained accomplishments on various tasks, such as the larger RegNets [91, 18] and Vision Transformers [92, 93, 94]. Also, ResNet variations have been tailored for scene recognition and can perform better than its general version [37].
Finally, CSAI is a challenging research topic on both experimental and psychological levels. However, society must face an essential subject and find intelligent solutions to stop its spread, speeding up police work and keeping more children safe.

### v2

@@ III-A Problem Definition
Any study with CSAI must consider the ethical concerns surrounding such data. Therefore, it is common to always design methods under a protocol not dissimilar to the Domain Generalization (DG) task.
@@
The remaining challenge is then to define which $\mathcal{S}^{train}$ domains can be used to train such a robust model, as well as the training procedure itself. As presented previously, our chosen classification task is Indoor Scene Classification. Whereas scene classification may be acknowledged as a homogeneous classification task [43], indoor and outdoor environments vary greatly in textures, objects, colors, shadows, framing, and many other features. Indoor environments contain large non-textured regions — rare in outdoor environments — making models that work well on outdoor images fail at indoor scenes [88]. Moreover, the distribution and variety of objects are much denser in indoor environments, making semantic segmentation more expensive and prone to error [67].
To solve Indoor Scene Classification with CSAI in mind, we propose to compare the performance of a classic transfer learning approach of fine-tuning a network pre-trained on ImageNet — our baseline — against a selection of self-supervised models known for generalizing to a plethora of downstream tasks — in some instances surpassing their supervised counterparts [19]. By comparing the results on our Indoor Scene Classification benchmark, we are able to select the best-performing model for this task, the code, and learned weights, which are then sent to law enforcement partners to do the final assessment on real CSAI.
@@ III-B Datasets
As mentioned in Section II-A, SSL is trained in two stages; the self-supervised stage trains a model using a pretext task, a pseudo-labeled classification and representation learning task meant to instill general representations into the model. This pre-trained model is then fine-tuned on a downstream task and evaluated. For each of these stages, (i) pretext, (ii) downstream, and (iii) evaluation, we have designed new sub-sets of existing public datasets, leaving (iv) the final CSAI evaluation and related dataset in the hands of partner law enforcement agents. Creating new sub-sets was necessary because most scene classification datasets with enough samples comprise both outdoor and indoor classes, and we wish to filter out the former. We present the selection for each stage next.
@@
We propose two combined datasets for the pretext self-supervised task: Indoors.all and Indoors.real. The former comprises all chosen indoor images from Hypersim [89], OpenRooms [90], InteriorNet [91], and Places365-Challenge [18], while the latter is then simply the Places365-Challenge indoor images. With this in mind, Indoors.all is called so as it contains not only real but also synthetic images and their rendered views (depth, segmentation, etc.). These views are used as pseudo-labels in the SSL pretext task, which we believe is more informative than random augmentations and could potentially produce better models, inspired by [67]. Table I summarizes the included datasets and their sizes.
TABLE I: Datasets used for the self-supervised pretext task. Size stands for the total number of images, and Synth stands for synthetic datasets. Combinations of the above public datasets.
@@
Indoors.all (ours) 7,778,152
Indoors.real (ours) 2,490,632
While most SSL models are readily available pre-trained on the well-known ImageNet dataset, we corroborate the finding from [18] that shows that models trained on object-centric datasets do not perform well on scene-centric tasks and vice-versa and even that training on both distributions may be beneficial. This comparison is further discussed with the experimental results in Section V.
@@
We introduce a new subset of Places — called Places8 — where classes are selected to highlight environments most common in CSAI. This is a smaller dataset than the ones used for the pretext task; it represents our downstream task and is used for fine-tuning the model post self-supervised learning.
@@
Class Test Train Val % Original Categories
bathroom 5,740 51,655 200 13.4 bathroom, shower
bedroom 11,112 100,012 600 25.9 bedchamber, bedroom, hotel room, berth, dorm room, youth hostel
child’s room 4,650 41,849 300 10.8 child’s room, nursery, playroom
classroom 3,751 33,763 200 8.7 classroom, kindergarden classroom
dressing room 2,432 21,889 200 5.7 closet, dressing room
living room 9,940 89,458 500 28.7 home theater, living room, recreation room, television room, waiting room
studio 1,404 12,633 100 3.3 television studio
swimming pool 1,505 13,547 200 3.5 jacuzzi, swimming pool
Total 40,534 364,806 2300 100
@@
Thus, we designed a small “custom dataset” from online images to check if the model performance is outside of the controlled nature of Places8. The dataset comprises 80 images, 10 images per class from the 8 Places8 classes: bathroom, bedroom, child’s room, classroom, dressing room, living room, studio, and swimming pool.
The OOD Scenes set is a sample of images taken from Google images, Bing images, and the Dollar Street dataset [92] in a 4:3:3 ratio. All images are free to share, modify, and use, including Dollar Street, licensed under CC-BY 4.0 Commercial.
@@
These sources were chosen to contrast the Places dataset data from the web with underrepresented data. For reproducibility, Places8 and OOD Scenes are available at Zenodo https://doi.org/10.5281/zenodo.13910526, upon request.
@@ IV Model Selection
In this section, we go through the model selection process. We recall that experiments that directly use CSAI are limited by law enforcement agents’ availability and local hardware requirements and are generally discouraged. For this reason, we perform model selection with data that approximates the problem, hence our creation of Places8 in partnership with agents (Section III-B). The following experiments then sought to best answer the three questions posed in Section III by using 5-fold cross-validation and 3×3$\times$3 × experiment repetitions, comparing model variants on Places8 while avoiding spurious results from training randomization.
@@ IV-A Experimental Setup
We conducted all experiments using 6 NVIDIA RTX GPUs A6000 with 48 GB each. We used PyTorch 1.10222https://pytorch.org as the deep learning framework with the Vision library for state-of-the-art self-supervised learning (VISSL) [^3] https://vissl.ai. ResNet-50 [93] was chosen as the backbone for all experiments. ImageNet pre-trained models are used directly from public “model zoos” while scene-centric SSL pre-trained models are trained from scratch or ImageNet-based weights; all models are later fine-tuned on Places8. The training times were optimized through mixed precision and activation checkpointing — allowing training to progress at 1.5 hours/epoch. The batch size was set to 1024 samples per GPU ($1024\times 6$). The LARS optimizer was used with learning rate scheduler of cosine half wave 3$\times$ restarts given the high batch size we chose, which follows the same training hyperparameters and guidelines from VISSL [94]. SSL pre-training was performed in 50 epochs, given our training budget, while fine-tuning was done on 100 epochs for Places8 with early stopping supported by the validation loss.
@@ IV-B SSL on Object-centric Data
Because training within an SSL protocol is expensive (Section IV-A), we sought to limit our exploration of this first question to a single SSL method and observe only the effect of adding object-centric learning to the mix; SwAV was the chosen method for this section for it is the best performing on the ImageNet classification downstream task of the four SSL methods (75.3% vs. 73.2% from second place Barlow Twins); we, therefore, study the 5 model variants yielded by questions Q1 and Q2 to isolate the impact of training first on ImageNet (Q1). The results can be found in Table III. This evaluation shows that significant improvements can be made using the widely available object-centric data to complement training, even when the downstream task is scene-centric. Comparing all experiments on all folds, we could measure with 97% confidence an average improvement of 3.1 percentage points in balanced accuracy on Places8 validation when starting with ImageNet SSL vs. models trained with scene-centric SSL from scratch.
@@ IV-C SSL on Scene-centric Data
When considering the inclusion of synthetic data, we compare SSL training with indoors.real versus indoors.all and found a 3% advantage to not using additional data (with 89.8% confidence). We hypothesize that the available synthetic data — while abundant in the number of samples — is limited in its diversity, with a large number of views being derived from the same 3D environment.
@@
TABLE IV: Comparing the impact of the inclusion scene-centric — real and synthetic — datasets for pre-training and the SSL training protocol. Results are reported as accuracy on the downstream Indoor Scene Classification task on the Places8 dataset.
@@ IV-D The Self-Supervised Learning Protocol
Our selected model is then comprised of the answers, following the Barlow Twins SSL protocol with the pipeline of object-centric SSL with ImageNet, scene-centric SSL with indoors.real, and fine-tuning on the Places8 downstream task. The final model was trained end-to-end with the full training set of Places8 (no folds) and evaluated on its test set. The results showed then 71.6% balanced accuracy. This is higher than what is commonly shown for the full Places dataset [43], but with fewer classes, it is impossible to do a fair comparison. We hope future studies consider using Places8, given its relationship to the CSAI final task.
@@ V Results on OOD Data and CSAI
Finally, in this section, we test our validated model — fine-tuned with the full training set of Places8 — on our two evaluation datasets. First, we consider the newly introduced OOD Scenes (Section III-B), where images were selected to represent a sample distribution distinct from the Places dataset.
Through our partnership with law enforcement agents, we test how our Indoor Scene Classification model performs on real CSAI data. This experiment on CSAI was executed in a GPU workstation owned by and located on the partner agency. No training is performed with either set.
@@ V-B CSAI
These experiments were conducted through our partnership with expert Brazilian Federal Police agents; our validated model weights and inference code were sent to the agents, who proceeded to run and collect metrics.
Because Indoor Scene Classification is a new task for CSAI data, the agents first selected 615 random samples from their allotted-for-research dataset. These samples were then annotated with scene labels. The randomness of the process showed that while a significant amount of the images were taken indoors (392), some were outdoors (118) or did not depict enough of the environment to allow for a label (105). The random selection showing most photos taken indoors corroborates the need for research into indoor scenes for CSAI. We follow with the 392 indoor scenes for our experiments.
Moreover, this subset is grouped into two categories of child abuse material: CSAI and Suspected CSAI. The CSAI category comprises the most serious and sensitive cases, while Suspected CSAI does not show any nudity yet contains hints of eroticism. Among these, 313 images are from the CSAI category and 302 from the Suspected CSAI category. We show the distribution of samples between classes and between these two CSAI categories in Fig. 4. Classes have different frequencies per CSAI category, with “studio” being especially prevalent in Suspected CSAI (the agent affirms there is a correlation between this kind of scene and erotic posing). In contrast, the scenes of “classroom” and “dressing room” are not seen in any image. While this is partly due to the random selection process, it hints at a lower prevalence of these scenes, at least in the analyzed data.
@@ VI Discussion
We have shown that combining publicly available self-supervision models with large scene classification datasets for a pretext task helps indoor classification on scenes prevalent in CSAI (but including synthetic data does not). However, the evaluation with real CSAI has highlighted that these datasets do not represent scenes recurrent in these sensitive images. Further work could analyze if adding people to the non-CSAI images affects the overall performance, as it could further introduce biases the model may leverage to drive classification.
In particular, the ambiguity among classes such as “bedroom”, “child’s room”, and “living room” seems to be hard for the model to distinguish, particularly in CSAI, an ambiguity corroborated by the presence of correlation shift [96] between the Places and CSAI data distributions. The shift is evident in a few examples: the model correlates the presence of people in the “dressing room” class and the presence of children in “child’s room”, but in CSAI, there are people in all images (by definition), and most of those people are children regardless of the place depicted.
This work applied self-supervision to indoor classification to aid CSAI triage and recognition, showing that SSL can leverage scene images to perform better than supervised learning. SSL performance was compared in multiple conditions, using real and synthetic images and public and sensitive data. The inference in real CSAI is vital, exposing the real distribution of scenes in this data and how a model trained with publicly available data performs. The results obtained should serve as a starting point for further investigations.
@@ VII Challenges and Future Work
The current solution can classify at most two of the scene classes with acceptable performance, which can be used for filtering specifically for bathrooms and bedrooms but not much else. Our analysis is, of course, limited to Indoor Scenes, one aspect of CSAI, and we know a general solution must combine this with other aspects such as object and body part detection and people; the — expected — appearance of the latter was shown to particularly influence the performance of a model trained on samples with few people depicted. Given our findings regarding the importance of scenery to CSAI, corroborated during the data labeling by the agents, we hope our first step to be an incentive for more research to be done within the Indoor Scene Classification field.
Additionally, all experiments used ResNet-50, which is fairly popular in SSL research, but multiple other architectures have gained much research interest after their sustained accomplishments on various tasks, such as the larger RegNets [25, 97] and Vision Transformers [98, 99, 100]. Also, ResNet variations have been tailored for scene classification and can perform better than its general version [43].
Finally, CSAI is a challenging research topic on both experimental and psychological levels. However, society must confront this essential subject and find intelligent solutions to stop its spread, speeding up police work and keeping more children safe.